OpenStudy (anonymous):

Find a cubic function with the given zeros. -7, 5, -3

OpenStudy (whpalmer4):

If you know all the zeros of a function f(x), you can create a polynomial by multiplying terms that = 0 at the zeros. For example, if the zeros were 1, 2, 3, the polynomial would be \[(x-1)(x-2)(x-3) = (x^2-3x + 2)(x-3) = x^3-6x^2+11x-6\]